# Pay for The Second-Best Service: A Game-Theoretic Approach Against Dishonest LLM Providers

### v2

@@ 1. Introduction
In recent years, large language models (LLMs) have demonstrated capabilities across a multitude of domains (Thirunavukarasu et al., 2023; Hong et al., 2024; Balunović et al., 2025). Following the release of ChatGPT (OpenAI, 2022), there has been a rapid and widespread adoption of LLMs among the general public, embedding them into a variety of commonplace activities. The inference processes of these models are predominantly controlled by companies, which provide access to individual users through interactive interfaces and Application Programming Interfaces (APIs). The growing user base has also triggered the development of a secondary market (e.g., TogetherAI (TogetherAI, 2023)), where providers aggregate resources to offer more affordable services.
This market expansion introduces a critical vulnerability: the potential for service provider dishonesty. Since users interact with LLMs in a “black-box” manner, a clear economic incentive exists for providers to substitute the advertised models with less capable alternatives to reduce their own operational costs. Such deceptive practices can manifest in various forms. For instance, when user is asking for a large model like Qwen3-235B-A22B (Yang et al., 2025), a provider might secretly switch to a smaller one, such as Qwen3-30B-A3B. It might also deploy a quantized version (e.g., using INT8 or FP8) (Gao et al., 2025). Additionally, providers might dynamically adjust a model’s reasoning effort (as in reasoning models like OpenAI o1(OpenAI, 2024a)) without user consent. The pay-per-token API billing also introduces new risks, allowing providers to inflate charges by inserting superfluous tokens that are difficult for users to detect (Kumar et al., 2025).
@@
Existing research has approached this challenge from a technical detection standpoint. Efforts include auditing model outputs for semantic behavioral drift (Chen et al., 2023; Eyuboglu et al., 2024) or token-level consistency (Miao and Fang, 2025; Sun et al., 2025a), and examining the inherent discrepancy across different models (Gao et al., 2025; Sun et al., 2025b; Cai et al., 2025). In this work, we tackle the problem through the lens of algorithmic game theory and computational economics. In other words, we propose a mechanism in which the providers would compete with each other, thus protecting user utility.
@@ 1.1. Our Contributions
We formulate a mechanism design problem in which providers determine their strategies based on the user’s delegation mechanism. We characterize the providers’ action space, which comprises cost and token sequence reporting. Our proposed model, which draws upon both mechanism design theory (Nisan and Ronen, 1999) and contract theory (Dütting et al., 2025), addresses a novel problem that has not been extensively studied in prior literature.
@@
We provide a theoretical guarantee under the continuous cost control model. Inspired by (Braverman et al., 2019) in the online learning domain, we first propose an impossibility theorem, proving that there is no $o(T)$-approximately incentive compatible mechanism that can guarantee a first-best user utility (formally defined later) asymptotically. Based on this, our objective is to propose a mechanism that guarantees a utility close to the second-best user utility.
@@ 2. Related Work
Numerous studies have investigated methods for auditing the disparities between the expected and actual outputs of commercial LLMs from various perspectives. Chen et al. (2023); Eyuboglu et al. (2024) investigates how LLMs change their behavior over time. Sun et al. (2025b) explores the idiosyncrasies present in the text generated by different language models, finding that these unique characteristics can persist even after the text has been rewritten, translated, or summarized by external LLMs. Sun et al. (2025a) studies the auditing of invisible reasoning tokens in reasoning LLMs, specifically examining the existence of token count inflation with low-effort token injection under conditions of increased partial embedding exposure and minimal token visibility. Cai et al. (2025) investigates the use of Trusted Execution Environments (TEEs), a hardware architecture, to provide cryptographic guarantees of model integrity. Gao et al. (2025) propose a concise and effective statistical testing scheme that quantifies changes in the output distribution due to quantization and watermarking by testing for the Maximum Mean Discrepancy between distributions.
@@
The advancement of large language models brings new economic scenarios and problems. Dütting et al. (2024) introduce the token auction model, a mechanism design framework that allows multiple LLM agents to influence generated content through bids. Bergemann et al. (2025) provide economic frameworks for token pricing, and Laufer et al. (2024); Sun et al. (2024) analyse the game-theoretic structure of fine-tuning and RLHF respectively. Saig et al. (2024) investigates how to incentivize LLM service providers to generate high-quality texts via cost-robust contracts.
@@
Algorithms from online learning can be utilized to control LLMs under uncertainty. Zhang et al. (2023) frame RLHF as a contextual bandit and use Thompson sampling to select the most informative comparison queries. Chen et al. (2024) study the exploration–exploitation trade-off when an LLM generates candidate answers with an $\epsilon$-greedy wrapper. Wang et al. (2025) treat safety auditing as an adversarial multi-armed bandit. Recent work (Bouneffouf and Feraud, 2025) systematically explore future directions for combining multi-armed bandits with LLMs. Drawing inspiration from mechanism design in (Braverman et al., 2019), we propose a novel model to address problems within the LLM scenario. This new context, in turn, presents new challenges for theoretical guarantees.
@@ 3.1. The User-Provider Delegation Game
    Cost control: For query $t$, the provider $d_{t}$ can secretly set $c_{t}\in[0,p_{d_{t}}]$ to a value other than the truthful one. For example, the provider could choose a cheaper, lower-quality model or sub-model to generate the response, or resort to quantization (Gong et al., 2024).
@@
    Token sequence reporting: For query $t$, the provider $d_{t}$ can report a token sequence $\tau^{\prime}_{t}$ that is longer than the actual generated sequence $\tau_{t}$ to inflate its payment. This can be done by adding inconsequential reasoning tokens, or by using techniques like watermarking (Zhao et al., 2024) to add tokens that are difficult for humans to discern. We assume such misreporting has negligible cost. [^2] Note that how such inflation is realized is out of scope for this work.
@@ 4. Second-Best Utility Guarantee Mechanism
Blind trust phase I.
@@
Blind trust phase II.
@@ 5.2. Main Results
$O(T^{1-\epsilon}\log T)$-dominant strategy of all model providers] If for all providers $i\in[K]$ and choices of cost $c\in[0,\eta_{i}p_{i}]$, the cost-performance functions $h_{i}(c),g_{i}(c)$ satisfies:
@@ 6.1. Setup
Our scenario involves three independent providers, each offering three models. When a user delegates a query to a provider, the provider selects a model according to its strategy, returns the results to the user, and reports the model and the number of tokens used (which may be inaccurate), thereby charging the user. In addition to our proposed optimal strategy, we consider several other strategies, as shown in Table 1. Provider 1 employs three models: gpt-5-high (gpt, 2025), deepseek-r1 (DeepSeek-AI et al., 2025), and gpt-5-medium. Provider 2 employs o3-mini (OpenAI, 2024b), o1-mini (OpenAI, 2024a), and deepseek-r1. Provider 3 offers claude-4.0 (cla, 2025), o1-mini, and deepseek-r1. All model pricing information is obtained from the official websites of the respective models.
@@
We select the maximum flow task from NLGraph (Wang et al., 2023) as the dataset for our experiments for the following reasons: (1) the maximum flow task enables the generation of a large number of experimental samples; (2) current LLMs’ performance on the task has not yet reached saturation. Following Wang et al. (2023), we use the “partial credit” as the evaluation metric for a single query (which is also the user’s reward), which ranges from 0 to 1, with higher values indicating better performance.
@@ Acknowledgements.
We would like to express our sincere gratitude to Kaifeng Lyu for providing valuable feedback on an early version of this manuscript. We are also thankful to Minrui Luo for carefully reading of our proofs and for providing several useful suggestions. Any remaining errors are our own.

### v3

@@ 1. Introduction
In recent years, large language models (LLMs) have demonstrated capabilities across a multitude of domains (thirunavukarasu2023large; hong2024metagpt; balunovic2025matharena). Following the release of ChatGPT (chatgpt2022), there has been a rapid and widespread adoption of LLMs among the general public, embedding them into a variety of commonplace activities. The inference processes of these models are predominantly controlled by companies, which provide access to individual users through interactive interfaces and Application Programming Interfaces (APIs). The growing user base has also triggered the development of a secondary market (e.g., TogetherAI (togetherai2023)), where providers aggregate resources to offer more affordable services.
This market expansion introduces a critical vulnerability: the potential for service provider dishonesty. Since users interact with LLMs in a “black-box” manner, a clear economic incentive exists for providers to substitute the advertised models with less capable alternatives to reduce their own operational costs. Such deceptive practices can manifest in various forms. For instance, when user is asking for a large model like Qwen3-235B-A22B (yang2025qwen3technicalreport), a provider might secretly switch to a smaller one, such as Qwen3-30B-A3B. It might also deploy a quantized version (e.g., using INT8 or FP8) (gao2025model). Additionally, providers might dynamically adjust a model’s reasoning effort (as in reasoning models like OpenAI o1(o12024)) without user consent. The pay-per-token API billing also introduces new risks, allowing providers to inflate charges by inserting superfluous tokens that are difficult for users to detect (kumar2025overthinkslowdownattacksreasoning).
@@
Existing research has approached this challenge from a technical detection standpoint. Efforts include auditing model outputs for semantic behavioral drift (chen2023chatgptsbehaviorchangingtime; 10.1145/3630106.3659047) or token-level consistency (miao-fang-2025-user; sun2025coincountinginvisiblereasoning), and examining the inherent discrepancy across different models (gao2025model; sun2025idiosyncrasieslargelanguagemodels; cai2025gettingpayforauditing). In this work, we tackle the problem through the lens of algorithmic game theory and computational economics. In other words, we propose a mechanism in which the providers would compete with each other, thus protecting user utility.
@@ 1.1. Our Contributions
We formulate a mechanism design problem in which providers determine their strategies based on the user’s delegation mechanism. We characterize the providers’ action space, which comprises cost and token sequence reporting. Our proposed model, which draws upon both mechanism design theory (nisan1999algorithmic) and contract theory (10.1561/0400000113), addresses a novel problem that has not been extensively studied in prior literature.
@@
We provide a theoretical guarantee under the continuous cost control model. Inspired by (braverman2019multi) in the online learning domain, we first propose an impossibility theorem, proving that there is no $o(T)$-approximately incentive compatible mechanism that can guarantee a first-best user utility (formally defined later) asymptotically. Based on this, our objective is to propose a mechanism that guarantees a utility close to the second-best user utility.
@@ 2. Related Work
Numerous studies have investigated methods for auditing the disparities between the expected and actual outputs of commercial LLMs from various perspectives. chen2023chatgptsbehaviorchangingtime; 10.1145/3630106.3659047 investigates how LLMs change their behavior over time. sun2025idiosyncrasieslargelanguagemodels explores the idiosyncrasies present in the text generated by different language models, finding that these unique characteristics can persist even after the text has been rewritten, translated, or summarized by external LLMs. sun2025coincountinginvisiblereasoning studies the auditing of invisible reasoning tokens in reasoning LLMs, specifically examining the existence of token count inflation with low-effort token injection under conditions of increased partial embedding exposure and minimal token visibility. cai2025gettingpayforauditing investigates the use of Trusted Execution Environments (TEEs), a hardware architecture, to provide cryptographic guarantees of model integrity. gao2025model propose a concise and effective statistical testing scheme that quantifies changes in the output distribution due to quantization and watermarking by testing for the Maximum Mean Discrepancy between distributions.
@@
The advancement of large language models brings new economic scenarios and problems. 10.1145/3589334.3645511 introduce the token auction model, a mechanism design framework that allows multiple LLM agents to influence generated content through bids. bergemann2025economicslargelanguagemodels provide economic frameworks for token pricing, and 10.1145/3589334.3645366; sun2024mechanismdesignllmfinetuning analyse the game-theoretic structure of fine-tuning and RLHF respectively. saig2024incentivizing investigates how to incentivize LLM service providers to generate high-quality texts via cost-robust contracts.
@@
Algorithms from online learning can be utilized to control LLMs under uncertainty. zhang2023banditrlhf frame RLHF as a contextual bandit and use Thompson sampling to select the most informative comparison queries. chen2024exploration study the exploration–exploitation trade-off when an LLM generates candidate answers with an $\epsilon$-greedy wrapper. wang2025auditingmab treat safety auditing as an adversarial multi-armed bandit. Recent work (bouneffouf2025surveymultiarmedbanditsmeet) systematically explore future directions for combining multi-armed bandits with LLMs. Drawing inspiration from mechanism design in (braverman2019multi), we propose a novel model to address problems within the LLM scenario. This new context, in turn, presents new challenges for theoretical guarantees.
@@ 3.1. The User-Provider Delegation Game
    Cost control: For query $t$, the provider $d_{t}$ can secretly set $c_{t}\in[0,p_{d_{t}}]$ to a value other than the truthful one. For example, the provider could choose a cheaper, lower-quality model or sub-model to generate the response, or resort to quantization (gong2024surveylowbitlargelanguage).
@@
    Token sequence reporting: For query $t$, the provider $d_{t}$ can report a token sequence $\tau^{\prime}_{t}$ that is longer than the actual generated sequence $\tau_{t}$ to inflate its payment. This can be done by adding inconsequential reasoning tokens, or by using techniques like watermarking (zhao2024sokwatermarkingaigeneratedcontent) to add tokens that are difficult for humans to discern. We assume such misreporting has negligible cost. [^2] Note that how such inflation is realized is out of scope for this work.
@@ 4. Second-Best Utility Guarantee Mechanism
Blind trust phase I
@@
Blind trust phase II
@@ 5.2. Main Results
$O(T^{1-\epsilon}\log T)$-dominant strategy of all model providers] If for all providers $i\in[K]$ and choices of cost $c\in[0,p_{i}]$, the cost-performance functions $h_{i}(c),g_{i}(c)$ satisfies:
@@ 6.1. Setup
Our scenario involves three independent providers, each offering three models. When a user delegates a query to a provider, the provider selects a model according to its strategy, returns the results to the user, and reports the model and the number of tokens used (which may be inaccurate), thereby charging the user. In addition to our proposed optimal strategy, we consider several other strategies, as shown in Table 1. Provider 1 employs three models: gpt-5-high (gpt-5), deepseek-r1 (deepseekai2025deepseekr1incentivizingreasoningcapability), and gpt-5-medium. Provider 2 employs o3-mini (o3mini), o1-mini (o12024), and deepseek-r1. Provider 3 offers claude-4.0 (claude), o1-mini, and deepseek-r1. All model pricing information is obtained from the official websites of the respective models.
@@
We select the maximum flow task from NLGraph (wang2023can) as the dataset for our experiments for the following reasons: (1) the maximum flow task enables the generation of a large number of experimental samples; (2) current LLMs’ performance on the task has not yet reached saturation. Following wang2023can, we use the “partial credit” as the evaluation metric for a single query (which is also the user’s reward), which ranges from 0 to 1, with higher values indicating better performance.
@@ Acknowledgements.
We would like to express our sincere gratitude to Kaifeng Lyu for providing valuable feedback on a preliminary version of this manuscript. We are also thankful to Minrui Luo for carefully reading of our proofs and for providing several useful suggestions. Additionally, we would like to thank Yu Chen, Zhaohua Chen, Yaoxin Ge, Zixin Gu, Xuanyu Li and Dengji Zhao for their helpful discussions in the early stages of this work. Any remaining errors are our own.